OpenStudy (anonymous):

If 20% of an object is submerged in water, what is its density?

OpenStudy (irishboy123):

a floating body displaces its own mass of the fluid in which it floats.....

OpenStudy (shamim):

The weight of the displaced water = weight of the object

OpenStudy (shamim):

But Weight w=mg

OpenStudy (shamim):

Here g is a constant

OpenStudy (shamim):

So mass of displaced water= mass of the object

OpenStudy (michele_laino):

hint: if I call with V_0 the volume of your object, then the volume of the part submerged is 0.2*V_0. Now we have to apply the Priciple of Archimewde, so we get: \[\Large {\delta _O} \times {V_0} \times g = {\delta _W} \times 0.2 \times {V_0} \times g\] where g is gravity, g=9.81 m/sec, and \[\Large {\delta _O},\;{\delta _W}\] are the densities of the object and the water respectively